# Efficient Nearest Neighbor based Uncertainty Estimation for Natural Language Processing Tasks

### v1

@@ 5 Experimental Settings
| Tasks | Datasets | Train | Dev | Test | |
@@
5.1 Tasks and Datasets
We measure the uncertainty estimation performance on Sentiment Analysis (SA), Natural Language Inference (NLI), and Named Entity Recognition (NER) in In-domain (ID) and Out-of-Domain settings. Dataset statistics are described in Table 1.
Sentiment Analysis (SA)
Natural Language Inference (NLI)
Named Entity Recognition (NER)
5.2 Existing Methods
Softmax Response (SR)
Temperature Scaling (TS)
is a calibration technique by which the logits (i.e., the outputs of the final layer before the softmax) are divided by a temperature parameter before applying the softmax function Guo et al. (2017). We optimized by L-BFGS on dev set loss.
Label Smoothing (LS)
Monte-Carlo Dropout (MC Dropout)
is a regularization technique that be used for estimating uncertainty in neural networks Gal and Ghahramani (2016). This approach involves executing the model times with varying dropout masks and computing the mean softmax output across these runs. In our experiments, we set for all evaluations, and the dropout rate is 0.1.
5.3 Training Setting
In all experiments, we train and evaluate the models on a single NVIDIA A100 GPU with 40GB of memory. We used 333microsoft/deberta-v3-base and 444microsoft/mdeberta-v3-base He et al. (2023), as the transformer encoder from Wolf et al. (2020) pre-trained model checkpoints. Cross-entropy loss is minimized by AdamW Loshchilov and Hutter (2019) with a linear scheduler Goyal et al. (2017). The batch size is 32, and gradient clipping is applied with the maximum norm of 1. The initial learning rate was set to 1e-5. All experiments are run five times, and we report the mean and standard deviation of the scores.
| Methods | IMDb (In-domain) | Yelp (Out-of-domain) | ||||||
| NN-UE | 0.950.12 | 9.021.39 | 98.640.12 | 7.970.61 | 1.450.15 | 4.171.52 | 98.230.39 | 9.920.61 |
| Methods | MNLI (In-domain) | SNLI (Out-of-domain) | ||||||
| NN-UE | 1.410.47 | 10.772.34 | 97.180.09 | 23.831.29 | 1.800.37 | 5.121.47 | 96.000.22 | 34.972.48 |
Datastore Construction
It is necessary to preserve the representation of the data for training a density estimator in Density Softmax and NN search in DAC and NN-UE. We maintain final layer representations corresponding to CLS tokens in SA and NLI. In NER, we stored the hidden representation of the final layer as a token representation corresponding to the beginning of the word.
-Nearest Neighbor Search
5.4 Evaluation Metrics
To evaluate the confidence calibration performance, we choose Expected Calibration Error (ECE) and Maximum Calibration Error (MCE). For selective prediction, we evaluate Area Under the Receiver Operator Characteristic curve (AUROC) and Excess-Area Under the Risk-Coverage curve (E-AURC). Evaluation metrics computation details are described in Appendix B.
6 Results
6.1 Sentiment Analysis
We present the uncertainty estimation results for sentiment analysis. Table 2 shows the results of in-domain and out-of-domain uncertainty estimation. NN-UE consistently outperforms existing methods in terms of ECE, MCE, and E-AURC. In AUROC, LS outperforms in OOD setting, but NN-UE outperforms existing methods in ID setting. Furthermore, the proposed method clearly outperforms DAC that uses an ensemble of neighbor search results for each hidden representation, by adding the label term.
6.2 Natural Language Inference
In NLI, we evaluate the uncertainty estimation performance (calibration and selective prediction) and the out-of-distribution detection performance.
6.2.1 Uncertainty Estimation
First, we show the results of in-distribution and out-of-distribution uncertainty estimation in the NLI task using the DeBERTaV3 model in Table 3. Similar to Section 6.1, NN-UE shows a significant improvement in ECE, MCE and E-AURC. Furthermore, in AUROC, NN-UE outperforms existing methods in both ID and OOD settings. Galil et al. (2023) have reported that improving calibration performance does not necessarily lead to improving selective prediction performance, but our proposed method improves both. On the other hand, the degree of improvement is greater for calibration performance. Specifically, the largest improvement is obtained on SNLI, where NN-UE reduces MCE by more than 31.49 % pt compared to SR.
| Methods | FPR@95 (↓) | AUROC (↑) | AUPR-In (↑) | AUPR-Out (↑) |
| SR | 93.752.68 | 52.200.91 | 68.110.26 | 35.762.03 |
| TS | 91.012.01 | 52.701.08 | 68.080.79 | 37.771.81 |
| LS | 94.213.09 | 52.361.08 | 68.500.54 | 35.382.17 |
| MC Dropout | 93.261.71 | 52.681.26 | 68.210.81 | 36.411.63 |
| Density Softmax | 93.362.24 | 51.220.78 | 67.220.24 | 35.461.93 |
| DAC | 92.951.63 | 52.781.45 | 68.511.01 | 36.381.94 |
| NN-UE (w/o label) | 91.331.21 | 53.220.74 | 68.510.63 | 37.961.18 |
| NN-UE | 88.491.37 | 54.142.01 | 69.021.07 | 40.222.19 |
6.2.2 Out-of-Distribution Detection
Following the previous study Tomani et al. (2023), we carried out the experiments in the out-of-distribution detection task. Out-of-distribution detection is the task that determines whether the data is in-domain or not. This task is based on the intuition that we want to return predictions with high confidence in ID but with low confidence in predictions in OOD. We evaluated the out-of-distribution detection performance by using maximum softmax probability as the uncertainty score, and report FPR@95 (the FPR when the TPR is 95%), AUROC, Area Under the Precision-Recall curve (AUPR)-in and AUPR-out. AUPR-in indicates the AUPR score when ID samples are treated as positive; AUPR-out is vice versa.
| Methods | (In-domain) | (Out-of-domain) | (Out-of-domain) | ||||||
| NN-UE (w/o label) | 3.370.71 | 33.153.65 | 17.630.66 | 8.780.62 | 24.911.81 | 70.104.03 | 14.610.67 | 35.267.16 | 65.418.11 |
| NN-UE | 1.780.32 | 26.0213.72 | 20.141.27 | 7.500.42 | 16.532.61 | 74.275.43 | 14.150.33 | 39.846.02 | 71.819.04 |
Table 4 shows the out-of-distribution detection results when using MNLI/SNLI datasets as ID/OOD, respectively, in model. NN-UE consistently shows the out-of-distribution detection performance improvement.
6.3 Named Entity Recognition
In order to evaluate NLP tasks other than simple multi-class classification, we evaluate the effectiveness of the proposed method for uncertainty estimation in NER. Since NER focuses on entities, it is necessary to obtain the confidence of the entity. In this research, we use the product of the confidence of the tokens that construct the entity as the confidence level of the entity.
Table 5 shows the results of in-distribution and out-of-distribution uncertainty estimation using the OntoNote 5.0 dataset in the mDeBERTaV3 model. The proposed method improves the uncertainty estimation performance in all cases except ECE of DAC in the and domains, often resulting in large improvements compared to the SR. Moreover, we can see that TS and MC Dropout, which are often effective in multi-class classification, are not effective in evaluating entities like NER. On the other hand, E-AURC in NER is consistently better without using the kNN-UE label term.
| Text | As long as you go into this movie with the understanding that it’s not going to contain any historical fact whatsoever, it’s not bad.<br /><br />It’s on par with Sam Raimi’s Ḧercules: The Legendary Journeys,̈ as far as plot, acting, humour, and production values are concerned. You’ll see the similarities at several points. Most of the fight scenes are not as good however and the film suffers from that. … |
| Label | negative |
| SR & NN-UE pred. | positive |
| SR conf. | 0.76 |
| NN-UE (w/o label) conf. | 0.71 |
| NN-UE conf. | 0.60 |
| 11 |
6.4 Case Study: Effects of the Label Term in NN-UE for a Misclassified Example
Table 6 shows SR and NN-UE confidences, and in NN-UE for a misclassified example. In this case, SR and NN-UE make incorrect prediction even though the true label is negative. However, the confidence is appropriately reduced by including the distances from the neighbors in NN-UE, compared to SR. Moreover, by using the information that there are only 11 examples in neighbors with the same label as the predicted label among the neighbors obtained by NN search, our NN-UE shows that the confidence is further reduced.
7 Analyzes
7.1 Impact of Top-
To understand the behavior of NN-UE, we evaluated the performance in uncertainty estimation when changing the number of neighbors during NN execution.
Figure 2 shows the results for SA, and Figure 3 shows the results for NER. As is noticeable in NER, the smaller , the better uncertainty estimation tends to be. Since our method averages the distance to the top examples, logits are scaled to be more limited to neighbors by reducing . It is assumed that the uncertainty estimation performance is slightly improved as the NN-UE scoring becomes more dependent on neighbor data if is small.
| Methods | SNLI | OntoNotes 5.0 |
| SR | 21.590.76 | 5.750.27 |
| TS | 21.640.07 | 5.790.17 |
| LS | 21.700.07 | 5.800.19 |
| MC Dropout | 396.861.10 | 101.980.83 |
| Density Softmax | 22.020.05 | 6.020.07 |
| DAC | 2346.6236.06 | 326.001.41 |
| NN-UE (w/o label) | 23.020.04 | 10.360.21 |
| NN-UE | 23.070.05 | 10.480.12 |
| OntoNotes 5.0 (In-domain) | OntoNotes 5.0 (Out-of-domain) | |||||||
| +PQ () | 1.960.31 | 31.3318.74 | 20.231.27 | 3.320.05 | 7.570.45 | 16.432.73 | 74.385.36 | 7.230.16 |
| +Clustering () | 1.920.31 | 28.5511.24 | 20.131.22 | 3.310.06 | 7.600.41 | 17.122.35 | 74.345.35 | 7.330.21 |
| +DR () | 2.140.37 | 33.5210.84 | 20.121.26 | 2.870.04 | 8.080.53 | 24.035.46 | 74.505.42 | 6.200.20 |
| Only DR () | 1.800.36 | 27.8513.80 | 20.131.29 | 3.410.10 | 7.540.45 | 16.422.73 | 74.305.44 | 7.750.24 |
| Methods | OntoNotes 5.0 | OntoNotes 5.0 |
| NN-UE | 100.0 | 100.0 |
| +PQ () | 21.30 | 51.68 |
| +Clustering () | 18.60 | 11.04 |
| +DR () | 0.02 | 0.04 |
| Only DR () | 43.98 | 20.35 |
7.2 Impact of Efficient Nearest Neighbor Search Techniques
In this subsection, we investigate the inference time and uncertainty estimation performance when applying approximate nearest neighbor search techniques and dimension reduction when executing NN search in NN-UE. As shown in Table 7, in the sequence labeling based NER that requires the NN search execution per token, it takes twice as much inference time as SR. On the other hand, in NN-LM Khandelwal et al. (2020), dimension reduction and approximate NN search techniques are effective to improve inference speed while maintaining perplexity He et al. (2021a); Xu et al. (2023). Therefore, inspired by these works for faster NN-LM, we investigate how the approximate nearest neighbor search techniques, such as Product Quantization Jégou et al. (2011) or clustering, and dimension reduction affect the uncertainty estimation and inference speed of our proposed method: NN-UE.
Product Quantization
Product Quantization (PQ) Jégou et al. (2011) is a data compression technique based on vector quantization. In PQ, a -dimensional representation is divided into subvectors and quantized by performing -means clustering on the vectors in each subspace. Vector quantization can significantly reduce the amount of memory occupied by vectors.555For example, raw datastore in kNN-UE is 636MB on OntoNotes 5.0 , but PQ reduces it to 10MB. In addition, by calculating the distance between compressed PQ codes, we can efficiently calculate the estimated value of the original Euclidean distance.
Clustering
The original NN-LM uses an inverted file index (IVF) technique that speeds up the search by dividing the representation into clusters by -means and searching for neighbors based on centroids. In this study, we evaluate the uncertainty estimation performance and inference speed when the number of clusters .
Dimension Reduction
In general, Transformer-based models such as PLM have high-dimensional token representations. In high-dimensional spaces, nearest neighbor search often suffer from the curse of dimensionality. To reduce this problem, we apply dimension reduction to NN-UE similar to He et al. (2021a). In this study, we use Principal Component Analysis (PCA) as a dimension reduction algorithm to reduce the dimension of the datastore representations and the query representation .
Results: Combination of PQ, Clustering, and Dimension Reduction
We evaluate the uncertainty estimation performance and inference speed when applying PQ, clustering, and dimension reduction are applied sequentially. The evaluations are performed on the OntoNotes 5.0 test set, and the results for different parameters of PQ, clustering and dimension reduction are shown in Appendix C. Table 8 shows the results on OntoNotes 5.0 and as ID/OOD, respectively. We can see that while the uncertainty performance is not significantly degraded when PQ and clustering are applied simultaneously to NN-UE, ECE and MCE are degraded when dimension reduction by PCA is further applied.666Distance recomputation does not mitigate this behavior, see Appendix D. On the other hand, E-AURC is not degraded when applying the combination of PQ, clustering, and dimension reduction. Also, we can obtain the improvement of inference speed rivaling the level of speed close to SR by applying PCA.
To deepen our understanding of the changes in the behavior of the uncertainty performance due to appling of approximate NN search techniques or dimension reduction in NN-UE, we calculated the coverage that how much the indices obtained when using the default exhaustive search are covered when applying PQ, clustering, and dimension reduction, sequentially. Table 9 shows the coverages on OntoNotes 5.0 and as ID/OOD settings, respectively.
We can see that while applying PQ and clustering simultaneously or applying PCA individually covers a certain percentage of the example indices obtained from NN-UE, which uses exhaustive search by default, applying PQ, clustering, and PCA simultaneously hardly covers any of the indices from the default NN-UE. It is assumed that applying PQ and PCA leads to coarse distance computation in a single subvector, which would correspondingly degrade the uncertainty estimation performance in NN-UE. On the other hand, if they are not applied simultaneously, the degradation in the uncertainty performance is limited.
8 Conclusion
In this paper, we proposed NN-UE, which estimates uncertainty by using the distance to neighbors and labels of neighbors. The experimental results showed that our method showed higher uncertainty estimation performance than existing uncertainty estimation methods in SA, NLI and NER. Furthermore, we showed that the uncertainty estimation performance can be further improved by adjusting the number of neighbors in NN-UE. Our method can greatly improve uncertainty estimation performance, especially in text classification tasks, with little degrading in inference speed. On the other hand, to address the degradation of the inference speed in token-level tasks such as NER, we investigated the effects of efficient neighbor search techniques in NN-UE. As a result, we found that product quantization, clustering, or dimension reduction improves inference speed without degrading the uncertainty estimation much more, unless combining all of them simultaneously.
9 Limitations
In this study, we focused only on the classification-based tasks. On the other hand, taking advantage of the recent growth of Large Language Models, uncertainty estimation in text generation is also attracting attention Fadeeva et al. (2023); Lin et al. (2024). Therefore, to investigate the effectiveness of NN-UE in text generation tasks is an interesting direction for future research. Furthermore, although NN-UE only used the representation of the last layer of the base model, exploring for an appropriate representation for uncertainty estimation is a future challenge.
Ethical Considerations
In this study, we used existing datasets that have cleared ethical issues following policies of published conferences. Therefore, they do not introduce any ethical problems. On the other hand, we have an ethical consideration about uncertainty estimation. Specifically, decision support systems with machine learning algorithms do not necessarily have a positive effect on performance. Jacobs et al. (2021) showed that collaboration with machine learning models does not significantly improve clinician’s treatment selection performance, and that performance is significantly degraded due to the presentation of incorrect recommendations. This problem is expected to remain even if uncertainty estimation methods are applied to machine learning models. In addition, introducing uncertainty estimation methods could conversely lead humans to give overconfidence in machine learning models, resulting in performance degradation.
Acknowledgements
Appendix A Training Settings for Density Estimator in Density Softmax
In Density Softmax Bui and Liu (2024a), we use RealNVP Dinh et al. (2017) which has two coupling structures. Table 10 shows the hyperparameters for training RealNVP as the density estimator in Density Softmax.

### v2

@@ 5 Experimental Settings
5.1 Tasks and Datasets
We measure the UE performance on Sentiment Analysis (SA), Natural Language Inference (NLI), and Named Entity Recognition (NER) in In-domain (ID) and Out-of-Domain (OOD) settings.555The datasets in SA and NLI were set up with reference to Xiao et al. (2022). Dataset statistics are described in Appendix A.
Sentiment Analysis (SA)
Natural Language Inference (NLI)
Named Entity Recognition (NER)
5.2 Existing Methods
We employ the simple baselines: Softmax Response (SR) Cordella et al. (1995), Temperature Scaling (TS) Guo et al. (2017), Label Smoothing Miller et al. (1996); Pereyra et al. (2017) and MC Dropout Gal and Ghahramani (2016). In addition, we use the recent strong baselines for UE: Spectral-Normalized Gaussian Process (SNGP) Liu et al. (2020), Posterior Networks (PN) Charpentier et al. (2020), Mahalanobis Distance with Spectral-Normalized Network (MDSN) Vazhentsev et al. (2022), E-NER Zhang et al. (2023), Density Softmax Bui and Liu (2024), and DAC Tomani et al. (2023). Details on baselines can be found in Appendix C. We have also experimented with a variant of NN-UE without the label term in Eq. 5, denoted by “w/o label” to emphasize the impact of the neighbor labels.
5.3 Training Settings
In all experiments, we train and evaluate the models on a single NVIDIA A100 GPU with 40GB of memory. We used 666 https://huggingface.co/microsoft/deberta-v3-base and 777https://huggingface.co/microsoft/mdeberta-v3-base He et al. (2023), as the Transformer encoder from Wolf et al. (2020) pre-trained model checkpoints. We use the cross-entropy loss in all experiments, including the optimization of hyperparameters in NN-UE. Batch size is 32, and the initial learning rate was set to 1e-5. The gradient clipping is applied with the maximum norm of 1. All experiments are run five times, and we report the mean and standard deviation of the scores.
Datastore Construction
It is necessary to maintain the representation of the data for training a density estimator in Density Softmax and NN search in DAC and NN-UE. We use the final layer representations corresponding to CLS tokens in SA and NLI. In NER, we stored the hidden representation of the final layer as a token representation corresponding to the beginning of the word.
-Nearest Neighbor Search
We use Douze et al. (2024) as the GPU-accelerated NN search toolkit. Unless otherwise specified, we fix the number of neighbors in NN search,888In Section 7.1, we conducted experiments to examine the behavior when varying over the set , with representing the median. and use which is an index for exact search in L2 norm, as the default in NN-UE.
| Methods | IMDb (In-domain) | Yelp (Out-of-domain) | ||||||
| SNGP | 4.180.30 | 22.694.83 | 98.530.15 | 9.951.17 | 4.890.59 | 21.284.68 | 98.100.27 | 11.422.14 |
| PN | 4.280.43 | 24.430.20 | 98.060.27 | 10.995.63 | 4.690.35 | 24.410.32 | 97.560.25 | 15.823.94 |
| MDSN | 4.450.43 | 23.975.05 | 98.480.08 | 10.250.86 | 5.320.92 | 21.332.91 | 98.000.20 | 11.123.53 |
| NN-UE | 0.950.12† | 9.021.39† | 98.640.12 | 7.970.61† | 1.450.15† | 4.171.52 | 98.230.39 | 9.920.61 |
| Methods | MNLI (In-domain) | SNLI (Out-of-domain) | ||||||
| SNGP | 10.450.56 | 35.4213.89 | 95.910.12 | 42.032.72 | 14.281.04 | 31.163.42 | 93.400.44 | 63.216.84 |
| PN | 33.830.51 | 37.100.71 | 96.960.10 | 26.331.22 | 32.010.61 | 35.370.58 | 95.570.29 | 40.944.49 |
| MDSN | 8.340.46 | 29.046.43 | 97.070.14 | 32.032.29 | 9.440.47 | 38.5913.94 | 96.110.12 | 38.913.06 |
| NN-UE | 1.410.47 | 10.772.34† | 97.180.09 | 23.831.29† | 1.800.37 | 5.121.47† | 96.000.22 | 34.972.48 |
5.4 Evaluation
To evaluate the confidence calibration performance, we choose Expected Calibration Error (ECE) and Maximum Calibration Error (MCE) Naeini et al. (2015). For selective prediction, we evaluate Area Under the Receiver Operator Characteristic curve (AUROC) and Excess-Area Under the Risk-Coverage curve (E-AURC) Geifman et al. (2019). Evaluation metrics computation details are described in Appendix D. In NER, we performed the evaluation of the UE performance with the flat recombination of the labels and the confidence for all tokens, respectively.
6 Results
6.1 Sentiment Analysis
In SA, we evaluate the confidence calibration, selective prediction and out-of-distribution detection performance.
Confidence Calibration and Selective Prediction
First, we present the UE results for sentiment analysis by differentiating the in-domain and out-of-main performance in Table 1. NN-UE consistently outperforms existing methods in terms of ECE, MCE, and E-AURC. In AUROC, LS outperforms in OOD setting, but NN-UE outperforms existing methods in ID setting. Furthermore, the proposed method clearly outperforms DAC that uses neighbor search results for each hidden representation with the additional label term. The lower UE performance than NN-UE in DAC is probably due to the difficulty in optimizing hyperparameters by comprising many layers.
Out-of-Distribution Detection
Following the previous study Tomani et al. (2023), we carried out the experiments in the out-of-distribution detection task, which determines whether a data instance is in-domain or not. This task is based on the intuition that we want to return predictions with high confidence in ID but with low confidence in predictions in OOD. We evaluated the out-of-distribution detection performance by using maximum softmax probability as the uncertainty score, and report FPR@95 (the FPR when the TPR is 95%), AUROC, Area Under the Precision-Recall curve (AUPR)-in and AUPR-out. AUPR-in indicates the AUPR score when ID samples are treated as positive; AUPR-out is vice versa.
Table 3 shows the out-of-distribution detection results when using IMDb/Yelp datasets as ID/OOD, respectively, in model. NN-UE consistently shows the out-of-distribution detection performance improvement.
| Methods | FPR@95 (↓) | AUROC (↑) | AUPR-In (↑) | AUPR-Out (↑) |
| SR | 82.519.49 | 63.185.14 | 69.512.57 | 54.708.48 |
| TS | 83.127.50 | 65.633.64 | 70.992.02 | 56.196.11 |
| LS | 86.884.27 | 62.172.83 | 69.501.51 | 51.383.81 |
| MC Dropout | 87.333.38 | 63.964.09 | 70.132.39 | 53.185.41 |
| SNGP | 81.923.46 | 63.273.07 | 68.832.10 | 55.913.20 |
| PN | 82.845.11 | 67.544.29 | 66.592.45 | 55.325.26 |
| Density Softmax | 87.543.14 | 58.734.33 | 67.342.57 | 49.194.36 |
| DAC | 84.984.19 | 64.656.18 | 70.693.59 | 54.817.29 |
| NN-UE (w/o label) | 75.872.16 | 70.441.70 | 74.771.44† | 63.392.24 |
| NN-UE | 73.555.01† | 71.112.92† | 73.802.19 | 65.013.45† |
6.2 Natural Language Inference
We show the results of in-domain and out-of-domain UE in NLI task using the DeBERTaV3 model in Table 2. Similar to Section 6.1, NN-UE shows the best UE performance, especially when the label term is included. Galil et al. (2023) have reported that improving calibration performance does not necessarily lead to the improved selective prediction performance, but our proposed method improves both types of metrics. On the other hand, the degree of improvement is larger for calibration performance. Specifically, the largest improvement is obtained on SNLI, where NN-UE reduces MCE by more than 31.49 % compared to SR. Additional experimental results on the Brier score can be found in Appendix E.
| Methods | (In-domain) | (Out-of-domain) | (Out-of-domain) | ||||||
| E-NER | 7.980.42 | 61.8727.06 | 19.441.81 | 17.420.88 | 40.465.33 | 74.324.47 | 25.422.09 | 59.1610.33 | 72.006.57 |
| NN-UE (w/o label) | 3.370.71 | 33.153.65 | 17.630.66† | 8.780.62 | 24.911.81 | 70.104.03 | 14.610.67 | 35.267.16† | 65.418.11 |
| NN-UE | 1.780.32 | 26.0213.72 | 20.141.27 | 7.500.42 | 16.532.61† | 74.275.43 | 14.150.33 | 39.846.02 | 71.819.04 |
6.3 Named Entity Recognition
To evaluate NLP tasks other than simple multi-class classification, we evaluate NN-UE in NER. Since NER focuses on entities, we use the product of the confidence of the tokens that construct a single entity as the confidence of the entity.
Table 4 shows the results of in-domain and out-of-domain UE using the OntoNote 5.0 dataset in . NN-UE shows the best performance in 4 cases, i.e., ECE or MCE, often resulting in large improvements over SR. On the other hand, E-AURC in NER is consistently better without using the NN-UE label term.999Label imbalance or large number of class can significantly affect E-AURC on NER when using kNN-UE with the label term. Details are in Appendix F. E-NER, a recent UE method specifically designed for NER, is close to NN-UE in its entity level selective prediction performance, but the calibration performance is not high.
NN-UE shows good UE performance even when the target domain is relatively far from source domain , such as . We have hypothesized that NN-UE might not work if the prediction target is too far from the training data distribution. If the prediction target is too far from the training data, the representation of the prediction from the model will be unreliable when compared to the prediction in the same domain as the training data. In general, methods based on feature distances assume that they maintain information relevant to the correctness of the prediction Postels et al. (2022). Our experiments have shown that the problem could be mitigated probably because the domains that the base models do not recognize are limited in the NLP community where there are many strong pre-trained models based on self-supervised learning such as DeBERTaV3.
| Scores | Correct Instances | Incorrect Instances | ||
| NN-UE (w/o label) | NN-UE | NN-UE (w/o label) | NN-UE | |
| 0.49 | 0.50 | 0.41 | 0.27 | |
| Confidence | 0.95 | 0.93 | 0.82 | 0.72 |
| Methods | MNLI | OntoNotes 5.0 |
| SR | 8.410.03 | 2.490.08 |
| TS | 8.420.07 | 2.510.08 |
| LS | 8.440.06 | 2.530.03 |
| MC Dropout | 157.520.51 | 39.810.39 |
| SNGP | 10.582.09 | - |
| PN | 9.110.07 | - |
| MDSN | 9.651.36 | - |
| E-NER | - | 2.510.12 |
| Density Softmax | 8.570.06 | 2.590.05 |
| DAC | 785.156.72 | 183.460.76 |
| NN-UE (w/o label) | 9.050.07 | 4.940.10 |
| NN-UE | 9.080.10 | 4.990.07 |
7 Analysis
7.1 Impact of Top-
To understand the behavior of NN-UE, we evaluated the performance in UE when changing the number of neighbors during NN execution.
Figure 3 and 4 show the results for SA and NER, respectively. As is noticeable in NER, the smaller , the better UE tends to be. These results suggest that our method requires that nearer examples to calibrate confidence, but more distant examples are not important. When calculating in Eq. 5, automatically adjusting the importance weights based on the order or distance of the retrieved nearest neighbors could further improve UE performance. Similar experimental and theoretical analysis of out-of-distribution detection using only NN distance also suggests that using -th example is preferable Sun et al. (2022). Providing a similar theoretical justification for our NN-UE is an interesting future direction.
| OntoNotes 5.0 (In-domain) | OntoNotes 5.0 (Out-of-domain) | |||||||
| + PQ | 1.960.31 | 31.3318.74 | 20.231.27 | 3.320.05 | 7.570.45 | 16.432.73 | 74.385.36 | 7.230.16 |
| + IVF | 1.920.31 | 28.5511.24 | 20.131.22 | 3.310.06 | 7.600.41 | 17.122.35 | 74.345.35 | 7.330.21 |
| + DR | 2.140.37 | 33.5210.84 | 20.121.26 | 2.870.04 | 8.080.53 | 24.035.46 | 74.505.42 | 6.200.20 |
| Methods | OntoNotes 5.0 | OntoNotes 5.0 |
| NN-UE | 100.0 | 100.0 |
| + PQ | 21.30 | 51.68 |
| + IVF | 18.60 | 11.04 |
| + DR | 0.02 | 0.04 |
| Only DR | 43.98 | 20.35 |
7.2 Importance of Label Term in
We analyze the impact of the label term Eq. 5 on the NN-UE confidence computation. We have shown that the UE performance is improved in several experiments. However, it is not obvious whether the improvement in UE performance is due to the reduction in primarily caused by adding the label term. Therefore, we examined the values for correctly and incorrectly predicted instances in both the absence and presence of the label term in NN-UE. Table 5 shows the distance terms, label terms and results for each case. If the predictions are correct, the growth of due to the label term is limited. On the other hand, NN-UE with label term remarkably reduce leading to the reduced confidence when the predictions are incorrect. This result suggests that the improvement of the evaluation metrics in NN-UE with label term is not achieved by increasing the confidence when the prediction is correct, but by appropriately reducing the confidence when the prediction is incorrect.
7.3 Impact of Efficient Nearest Neighbor Search Techniques
We investigate the inference time and UE performance when applying approximate nearest neighbor search techniques and dimension reduction when executing NN search in NN-UE as a real world application. As shown in Table 6,101010Other results can be found in Appendix G. in the sequence labeling based NER, which requires executing NN searches per token, it takes twice as much inference time as SR.111111Inference times do not increase as dramatically as -Nearest Neighbor Language Model Khandelwal et al. (2020) because NN can be executed in parallel for both classification and NER. On the other hand, in -Nearest Neighbor Language Model (NN-LM) Khandelwal et al. (2020), dimension reduction and approximate NN search techniques are effective to improve inference speed while maintaining perplexity in text generation He et al. (2021a); Xu et al. (2023). Therefore, inspired by these works for faster NN-LM, we investigate how the approximate nearest neighbor search techniques, such as Product Quantization Jégou et al. (2011) (PQ), Inverted File (IVF) clustering and dimension reduction affect the UE and inference speed of our NN-UE. Description of approximate nearest neighbor search techniques and detailed discussion when each method is individually applied to NN-UE are in Appendix I.
Results of Combination of PQ, IVF and Dimension Reduction
We evaluate the UE performance and inference speed when applying PQ, IVF and dimension reduction are applied. Table 7 shows the results on OntoNotes 5.0 and test sets as ID/OOD, respectively. The detailed discussion when changing the parameters of PQ, clustering and dimension reduction are shown in Appendix I. We can see that ECE and MCE are degraded when PQ, IVF and dimension reduction by PCA are applied simultaneously to NN-UE.121212Distance recomputation does not mitigate this behavior, see Appendix J. On the other hand, our results show that applying them appropriately such as combining PQ with IVF improve inference time with mitigating the degradation in UE performance (The results with the parameters for PQ or IVF can be found in Appendix I.1 or I.2). To deepen our understanding of the above changes in the behavior of the uncertainty performance due to appling of approximate NN search techniques or dimension reduction in NN-UE, we calculated the coverage that how much the indices obtained when using the default exhaustive search are covered when applying PQ, clustering, and dimension reduction sequentially.
Table 8 shows the coverages on OntoNotes 5.0 and as ID/OOD settings, respectively. We can see that applying PQ, clustering, and PCA simultaneously hardly covers any of the indices from the default NN-UE. It is assumed that applying PQ and PCA in the same time leads to coarse distance computation in a single subvector, which would correspondingly degrade the UE performance in NN-UE. Actually, the experimental results in Table 17 in Appendix I.3 suggest that excessive dimension reduction in distance computation could have a negative impact on the UE performance. On the other hand, if combined with PQ and IVF, or applied PCA individually, some of the ground-truth nearest neighbor examples still exist.
8 Conclusion
In this paper, we proposed NN-UE, which estimates uncertainty by using the distance to neighbors and labels of neighbors. The experimental results showed that our method showed higher UE performance than existing UE methods in SA, NLI and NER. Furthermore, our analysis of correctly and incorrectly predicted instances suggests that the improvement in NN-UE is largely due to the reduction in confidence on incorrect instances. In addition, we investigated the effects of efficient neighbor search techniques in NN-UE to address the degradation of the inference speed in token-level tasks such as NER. As a result, we found that product quantization, clustering, or dimension reduction improves inference speed without degrading the UE much more, unless combining all of them simultaneously.
9 Limitations
In this study, we focused only on the classification-based tasks. On the other hand, taking advantage of the recent growth of Large Language Models, UE in text generation is also attracting attention Yoshikawa and Okazaki (2023); Fadeeva et al. (2023); Lin et al. (2024). Therefore, to investigate the effectiveness of NN-UE in text generation tasks is an interesting direction for future research. Not only that, our proposed method is applicable to more tasks such as image classification.
Furthermore, although NN-UE only used the representation of the last layer of the base model, exploring for an appropriate representation for UE is a future challenge. Also, to investigate the relationship between the representation quality and in- and out-of-domain UE performance when using smaller pretrained encoders than DeBERTa, such as BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019) is an interesting direction.
Finally, we used ECE and MCE to measure calibration performance. On the other hand, it may be more appropriate to use other metrics to measure calibration performance when the dataset with multiple annotations including human disagreement is available Baan et al. (2022), where it may be similar to the label disagreement in similar output representations. In Section 7.2, we showed that our NN-UE with the label term improves in the direction we expected: it reduces confidence much more when the predictions are inaccurate. However, measuring calibration performance on a variety of data with multiple annotations may provide a more interesting insight into the behavior of our proposed method.
Ethical Considerations
In this study, we used existing datasets that have cleared ethical issues following policies of published conferences. Therefore, they do not introduce any ethical problems. On the other hand, we have an ethical consideration about UE. Specifically, decision support systems with machine learning algorithms do not necessarily have a positive effect on performance. Jacobs et al. (2021) showed that collaboration with machine learning models does not significantly improve clinician’s treatment selection performance, and that performance is significantly degraded due to the presentation of incorrect recommendations. This problem is expected to remain even if UE methods are applied to machine learning models. In addition, introducing UE methods could conversely lead humans to give overconfidence in machine learning models, resulting in performance degradation.
Acknowledgements
Appendix A Dataset Statistics
The dataset statistics in our study is shown in Table 9.
| Tasks | Datasets | Train | Val | Test | |
@@ Appendix A Dataset Statistics
Appendix B Training Settings for Density Estimator in Density Softmax
In Density Softmax Bui and Liu (2024), we use RealNVP Dinh et al. (2017) as the density estimator, which has two coupling structures. Table 10 shows the hyperparameters for training RealNVP as the density estimator in Density Softmax.